If Meena gives an interest of ₹ 45 for one year at 9% rate p.a. What is the sum she has borrowed?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that, SI =₹45, R = 9%, T = 1 year, P =?
SI = (P × R × T)/100
45 = (P × 9 × 1)/ 100
P = (45 ×100)/ 9
= 5 × 100
=₹500
Hence, she borrowed₹500.

Thanks